# Accelerating AllReduce with a Persistent Straggler

### v1

@@ 4.2 Evaluating StragglAR on Multi-GPU Servers
Figure 5: Results on 8 A100 GPUs with NVSwitched connectivity. (a) shows the optimistic use case, by assuming ReduceScatter can complete within the straggler delay. (b) uses a straggler delay of 9.45 ms, the average from §2. (c) fixes the buffer size at 4 GiB and varies the straggler delay.
@@
We run experiments on different multi-GPU setups. Our primary testbed is a VM with 8 80GB Nvidia A100 GPUs connected by NVSwitch [17], enabling any-to-any connectivity with 600 GB/s bidirectional P2P bandwidth. We also run experiments on one node of the Perlmutter supercomputer [16] with 4 40GB Nvidia A100 GPUs fully connected with NVLink 3.0, enabling 200 GB/s bidirectional P2P bandwidth (results in §D). We use CUDA events on GPU rank 0 to capture the runtime. End-to-end training experiments can be found in §F and §G.
@@
Varying buffer size. We first assume the straggler delay can mask the ReduceScatter precondition and measure runtime from this point to capture performance for ideal use cases of our algorithm. (For Broadcast, we assume the entire AllReduce precondition has completed.) Following the standard for benchmarking AllReduce, we evaluate the communication time as we scale the buffer size in powers of 2 (in bytes). NCCL behaves unpredictably when the P2P data transfer size (chunk size) is not a multiple of 1 KiB, so we pad buffers for StragglAR to ensure the chunk size is the lowest multiple of 1 KiB greater than $\tfrac{s}{n-1}$ (for baselines, the chunking scheme inherently ensures this when $s$ is a power of 2). For each buffer size, we run 50 iterations per algorithm and report the mean, with error bars for the standard error of the mean.
@@
Similar to how NCCL reports collective performance [31], we show the algorithmic bandwidth—the buffer size divided by the collective communication time (i.e., runtime normalized for input size)—in Fig. 5a. For some smaller buffer sizes, RHD and Broadcast outperform StragglAR, which still outperforms Ring; this matches results expected from $\alpha{-}\beta$ costs since $\alpha$ costs dominate with smaller buffer sizes. However, StragglAR is consistently the fastest for the larger buffer sizes (which scale with model size), specifically 8.3% and 22% faster than Ring on the 4- and 8-GPU servers, respectively. The outlier for 256 MiB is likely caused by unexpected chunking behavior implemented internally by NCCL, as shown from running the standard nccl-tests profiler (see §E).
@@
To simulate a persistent straggler, we idle GPU rank $n{-}1$ for the specified number of clock cycles. Simultaneously, we begin the ReduceScatter for the other $n{-}1$ GPUs for StragglAR (similarly, AllReduce for Broadcast), thereby starting at the $T_{rest}$ point described in §2. We first conduct the same experiments as before, but instead assume the average delay of 9.45 ms observed in our profiling experiments in §2. Fig. 5b shows that StragglAR improves the runtime under typical straggler delays in distributed ML with larger buffer sizes, with average-case performance typically matching ideal performance. Interestingly, for the 4 GiB buffer, StragglAR’s performance declines slightly from the ideal case because of the longer, non-overlapped ReduceScatter precondition that diminishes performance improvements from our straggler-aware schedule.
@@
Varying straggler delay. We systematically varied the straggler delay by adjusting the number of clock cycles for the sleep kernel, and ran all algorithms on the largest buffer size previously tested (1 GiB for the 4-GPU setup and 4 GiB for the 8-GPU setup) as a stress test of the precondition. Figs. 5c and 7c capture the total time from when the AllReduce kernel is called on non-straggler GPUs to when it completes. The results show the critical delays of 7.15 ms (8-GPU, 4 GiB) and 4.11 ms (4-GPU, 1 GiB) after which StragglAR outperforms baselines in each setup. However, this critical delay depends on the buffer size, as shown in Fig. 6(a), which plots the runtime of the ReduceScatter precondition on the VM with 8 A100 GPUs. While the shaded region in Fig. 6(a) captures the straggler delays at which StragglAR achieves its full theoretical gains, StragglAR still outperforms baselines when some, but not all, of the ReduceScatter can be overlapped with the straggler delay. The critical delay with a 4 GiB buffer in Fig. 5c is almost 2×2$\times$2 × less than the ReduceScatter runtime for 4 GiB. This indicates that StragglAR enables significant speedups even when the precondition is not fully masked. Specifically, if the straggler delay is shorter than the ReduceScatter, StragglAR still outperforms baselines as long as $T_{delay}\geq T_{RS}-\max\{T_{B}-T_{SAR},0\}$, where $T_{RS}$ is the ReduceScatter runtime, $T_{delay}$ is the delay, and $T_{B}$ is baseline $B$’s runtime.
@@ 4.3 Scaling characteristics of StragglAR
(a) Ideal straggler delay.
(b) Speedup over Ring AllReduce as cluster size scales.
Figure 6: (a) shows the ReduceScatter runtime over buffer sizes on the 8-GPU RunPod server to highlight a possible range of straggler delays when StragglAR achieves full theoretical guarantees. (b) compares the AllReduce scaling behavior with the analytical network model for 1 GiB buffer.
Because we lack access to the largest scale-up domains available today, we simulate performance on clusters of up to 256 GPUs—–the maximum supported by current hardware—–by using the analytical network model, similar to prior work [21, 32]. The analytical model uses the $\alpha{-}\beta$ costs to simulate the AllReduce performance. We use $3$ $\mu s$ for $\alpha$ based on latency profiling studies of recent NVLinks [33] and the inverse of 450 GB/s, the NVLink unidirectional bandwidth on H100 SXMs, for $\beta$. In these simulations, we assume sufficient straggler delay to mask the ReduceScatter precondition for StragglAR (and the AllReduce precondition for Broadcast), which is the typical use case for StragglAR. We find that StragglAR outperforms all baselines for realistic, large buffer sizes (1 GiB) both when the number of GPUs is a power-of-2 and a non-power-of-2 (by evaluating at exactly the midpoint value of $n$ between each power of 2), as shown in Fig. 6. At $n=256$, StragglAR provides nearly 2×2$\times$2 × speedup over the Ring algorithm in settings with a persistent straggler.
Limitations. Firstly, StragglAR primarily provides benefits for workloads with persistent stragglers that typically exceed the critical delay. If performance degrades, the user can re-profile the workload to determine if the persistent straggler has changed and modify the runtime accordingly (by swapping straggler ranks once offline). Second, the critical delay may vary depending on the hardware setup because the GPU model, CUDA/NCCL versions, NVLink bandwidth, etc., impact performance. Further, the algorithm does not support odd values of $n$, though this is an atypical setup for large-scale distributed ML. Finally, although StragglAR can benefit workloads with multiple stragglers, since a “straggler” is by definition relative, it is less applicable when several GPUs simultaneously straggle together; however, this scenario would be uncommon due to tail effects (§2[3].
@@ 5 Conclusion
We design StragglAR, an AllReduce algorithm for faster, bandwidth-efficient AllReduce in distributed ML workloads with a persistent straggler. StragglAR provides a 2×2$\times$2 × theoretical speedup over the Ring algorithm in large clusters, and a 22% speedup in practice on an 8-GPU server.
@@ Acknowledgements
Appendix A Related Work
@@
Optimal collective communication. Over decades, researchers have implemented collective communication primitives with fast algorithms that are carefully attuned to latency and bandwidth trade-offs. With today’s large ML models that require communicating large amounts of data, bandwidth-efficient algorithms are of primary focus [34, 35]. In homogeneous networks, well-known parallel algorithms, such as Ring, recursive halving/doubling, etc., enable fast, bandwidth-efficient AllReduce by heavily parallelizing communication in every round of the algorithm [30, 36, 37, 19, 28, 38]. For heterogeneous networks, recent works synthesize custom collective algorithms by using the hardware topology or parallelization strategy as an input [20, 39, 40, 41, 42, 43, 44]; optimizing a collective algorithm to minimize the latency and bandwidth cost, however, is a challenging combinatorial optimization problem that often requires hand-tuned heuristics to constrain the search space [20]. Other works provide flexible and hardware-aware communication primitives for designing and optimizing collective communication algorithms [45, 46]. Further, some approaches optimize AllReduce for a specific topology [47, 48, 49]. Our work extends this line of research by optimizing AllReduce in settings with a persistent straggler and when data buffers are large and thus require bandwidth-efficient algorithms, such as in distributed ML [34, 20, 50].
@@
Rebalancing workloads. Recent works improve ML training and inference efficiency by optimizing the model parallelization strategy, which can affect data/work sharding across GPUs and the collective communication incurred [51, 52, 53, 54, 55, 10, 34, 56]. Vanilla parallelization strategies evenly shard or replicate the compute and memory load (i.e., model weights, input batches, activations, optimizer states, etc.) across GPUs within a group to achieve speedups from parallelization. This balances the computation and communication load experienced by each worker, but could introduce resource idling when there are slow-running GPUs in the group. Our experiments in §2 highlighted a significant and persistent straggler even with uniform workload placement. Other works identify severe stragglers and balance the workload dynamically to limit their impact [4, 1, 57, 58]. StragglAR directly addresses stragglers at a lower layer of abstraction, within the AllReduce kernel, and thus expands on the problems tackled by these works.
@@
Sending less data. Various methods leverage the resilience of stochastic optimization algorithms used in training to reduce the communication volume by selectively communicating data with high statistical value [3, 59, 60], synchronizing gradients with varying frequencies [61, 62, 63, 5, 64, 64], or sending lower-precision data [65, 66, 67, 68, 60, 69, 70].  StragglAR is orthogonal to these works because it optimizes the collective algorithm itself (i.e., exactly how the data is transmitted as opposed to how much) and can operate given any buffer size, including after compression. Further, because StragglAR directly modifies the collective algorithm, as opposed to the data that is sent or the frequency of synchronization, it also supports AllReduce for tensor parallelism, which is used in both training and inference [7].
Asynchronous collectives. Asynchronous paradigms are emerging as an alternative to prevent synchronization barriers from delaying collective communication [71, 72, 73, 74, 75, 76, 77, 78, 79, 80]. While these schemes are used for smaller, irregular communication patterns, such as expert-parallel AllToAll [81], they are less useful with large data buffers (such as in distributed ML AllReduce) that cause network bandwidth to become the bottleneck [19]. Thus, bulk data transfers, like gradient synchronization in data-parallel training, still requires using collective algorithms that minimize communication time by efficiently using bandwidth.
Appendix B Proof of Communication Complexity

### v2

@@ 4.1 Benchmarking StragglAR on Multi-GPU Scale-Up Domains
Figure 5: Benchmarking AllReduce performance for different algorithms on NVIDIA DGX H100 (a-c) and A100 (d-f) servers. (a), (d) show the optimistic use case, by assuming ReduceScatter can complete within the straggler delay. (b), (e) use a straggler delay of 9.45 ms, the average from the ML workloads in Fig. 2(a). (c), (f) fix the buffer size at 4 GiB and vary the straggler delay.
@@
Setup. We run experiments on several multi-GPU testbeds: (1) NVIDIA DGX H100, $\times8$0GB H100 GPUs connected by NVLink 4.0 (450 GB/s P2P) and NVSwitch 3.0; (2) NVIDIA DGX A100, $\times8$0GB A100 GPUs connected by NVLink 3.0 (300 GB/s P2P) and NVSwitch 2.0; (3) a node of the Perlmutter supercomputer [NERSC, 2025], with $\times4$0GB A100 GPUs fully connected (mesh) by NVLink 3.0 for 100 GB/s P2P (results in §F). (1) and (2) are VMs obtained through RunPod [RunPod, Inc., 2025]. We use CUDA events on GPU rank 0 to measure the runtime.
@@
Varying buffer size. We first assume the straggler delay can mask the ReduceScatter precondition and measure runtime from this point to capture performance for ideal use cases of our algorithm. (For Broadcast, we assume the entire AllReduce precondition has completed.) Following the standard for benchmarking AllReduce performance, we evaluate the communication time as we scale the buffer size in powers of 2 (in bytes). NCCL behaves unpredictably when the P2P data transfer size (i.e., chunk size) is not a multiple of 4 KiB (page size), so we pad buffers for StragglAR to ensure the chunk size is the lowest multiple of 4 KiB greater than $\tfrac{s}{n-1}$ (for baselines, the chunking scheme inherently ensures this when $s$ is a power of 2). For each buffer size, we run 50 iterations per algorithm and report the mean, with error bars for the standard error of the mean.
@@
Similar to how NCCL reports performance [NVIDIA, 2024b], we show the algorithmic bandwidth—buffer size divided by collective communication time (i.e., runtime normalized by input size)—in Fig. 5 (a,d). For some smaller buffer sizes, RHD, Broadcast, and MSCCL outperform StragglAR, which still outperforms Ring; this matches expected results from $\alpha{-}\beta$ costs since $\alpha$ costs dominate for small buffers. However, StragglAR is consistently the fastest for the larger buffer sizes (which scale with model size), specifically 8.3% faster on the 4-A100 Perlmutter node and >25% faster on the 8-H100 and 8-A100 DGX servers. The outlier at 256 MiB likely stems from NCCL’s internal tuning in the 64-512 MiB range, where internal protocol changes cause unexpected performance, as confirmed by our own nccl-tests profiling (§G) and prior work [Xu et al., 2025, Hu et al., 2025].
@@
To simulate stragglers, we idle GPU rank n1n{-}1 for the specified number of clock cycles. Simultaneously, we begin the ReduceScatter for the other n1n{-}1 GPUs for StragglAR (similarly, AllReduce for Broadcast). We conduct similar experiments as above, but instead impose the average delay of 9.45 ms observed in our Llama-3.2 fine-tuning experiments in Fig. 7. Fig. 5 (b,e) shows that StragglAR improves the runtime under typical straggler delays in distributed ML with larger buffer sizes, with average-case performance closely matching ideal performance. For the 4 GiB buffer, StragglAR’s performance declines slightly from the ideal case because the ReduceScatter for this buffer size cannot be fully overlapped with the 9.45 ms straggler delay.
@@
Varying straggler delay. We vary the straggler delay by adjusting the number of clock cycles for the sleep kernel, and run all algorithms on the largest buffer size (4 GiB) as a stress-test of the precondition overlap. Fig. 5 (c,f) captures the total time from when the AllReduce is called on non-straggler GPUs to when it completes. It shows the critical delays of 5.53 ms and 7.57 ms for the DGX H100 and A100, respectively, after which StragglAR outperforms baselines. The critical delay is higher for the DGX A100 since a ReduceScatter takes longer due to its lower P2P bandwidth. The critical delay also depends on buffer size, as shown in Fig. 6(a), which plots the runtime of the ReduceScatter precondition on the DGX H100. While the shaded region in Fig. 6(a) captures the straggler delays at which StragglAR achieves its full theoretical gains, StragglAR still outperforms baselines when some, but not all, of the ReduceScatter is overlapped with the straggler delay. For example, the critical delay for a 4 GiB buffer in Fig. 5c is less than the ReduceScatter time for 4 GiB. Thus, StragglAR enables speedups even if the straggler delay is less than the ReduceScatter precondition, but longer than the critical delay. As we show in §4.3, the critical delay approaches zero as the size of the GPU cluster increases due to StragglAR’s efficient scaling.
4.2 End-to-end Evaluation on ML Workloads
We run experiments on DGX A100 VMs to assess StragglAR’s impact on end-to-end training time with data-parallel fine-tuning of three popular LLMs: Llama-3.2-3B [Grattafiori et al., 2024], Phi-3-mini-3.8B [Abdin et al., 2024], and Qwen-2.5-3B [Qwen, 2024]. We compile our AllReduce implementations into a custom package and configure PyTorch’s AllReduce to call this backend. In each VM, we first profile the workload with standard PyTorch tools and identify persistent stragglers, i.e., ranks that are most likely to be delayed based on data from prior runs. We use this approach both because persistent stragglers are common according to prior work [Lin et al., 2025, Jiang et al., 2024] and our own experiments (§B), and because it stress-tests StragglAR to encounter both ideal and worst-case conditions (when a different/no rank is the straggler). Then, we pass the selected rank to the backend and train the model for 100 iterations (batch size of 32). Because the buffer sizes of these models are $\sim$4 GiB, we compare StragglAR to the Ring algorithm, which is optimal at large buffer sizes (Fig. 5a). We show the end-to-end speedups in Table 2.
Model Speedup Straggler GPU hrs.
(%) persistence (%) saved/day
Llama-3.2-3B 4.75 90 9.12
Phi-3-mini-3.8B 4.43 95 8.51
Qwen-2.5-3B 2.39 77 4.59
Table 2: End-to-end training speedups over Ring, and GPU-hours saved per day on an 8-GPU server.
The end-to-end speedup depends on several factors: how often the actual straggler was the rank we passed to the backend, the straggler’s delay, and the fraction of overall time spent on AllReduce. For Qwen-2.5-3B, gains are smaller because the straggler on that VM was less persistent, leading our algorithm to encounter its worst-case scenario (i.e., no ReduceScatter overlap) more often. Even then, StragglAR consistently shows end-to-end gains, as its worst-case performance nearly matches baselines while its upside is much higher (Fig. 2(b)). StragglAR does not require online straggler detection with dynamic stragglers, as eager conditional execution of schedules based on the first n1n{-}1 ready ranks means at worst (i.e., no straggler delay), StragglAR’s performance closely matches baselines. These speedups translate to 9.12 GPU-hours saved per day on an 8-GPU server (Table 2).
@@ 4.3 Scaling Characteristics of StragglAR
(a) Ideal straggler delay (DGX H100).
(b) Speedups over Ring as cluster size scales.
Figure 6: (a) ReduceScatter time over buffer sizes, capturing the range of straggler delays where StragglAR achieves full theoretical guarantees. (b) AllReduce performance scaling with the analytical model for a 1 GiB buffer. The shaded region is the range of StragglAR’s possible speedups.
[…]
Limitations. To consistently achieve its best-case performance with dynamic stragglers, StragglAR requires conditional execution of schedules based on the n1n{-}1 ranks that are first ready, which can be complex. We note that libraries like NCCL already support conditional execution of different algorithms and protocols based on runtime conditions. While StragglAR matches baselines at scale even without a straggler, its performance on smaller clusters depends on the critical delay, which varies with the exact GPU P2P bandwidth. Our algorithm also does not support odd values of $n$, though such setups are atypical in large-scale ML. Finally, while StragglAR can tolerate multiple stragglers, since a straggler is by definition relative, it is less effective when many GPUs straggle simultaneously; however, this scenario is highly improbable since GPU execution times are continuous variables [Dean and Barroso, 2013, Warraich et al., 2025].
@@ 5 Conclusion
We design StragglAR, a parallel algorithm that exploits natural variation in GPU execution times to speed up AllReduce. StragglAR achieves a $2\times$ speedup over the known lower bound for bandwidth-optimal AllReduce, while performing as well in the worst case. By introducing the dimension of temporal asymmetry, StragglAR offers a new paradigm for collective algorithm design.
@@ Acknowledgements
Appendix
Appendix A Extended Related Work
We provide more discussion of related work in this section.
In distributed training and inference, each GPU produces local gradients and activations that must be aggregated across devices. This multi-device aggregation is performed using the AllReduce collective communication primitive, which transmits and reduces data across the inter-GPU network using optimized parallel communication algorithms. The parallelization strategy of an ML job determines which GPUs communicate and how often: data parallelism uses AllReduce regularly to average gradients, while tensor model parallelism [Shoeybi et al., 2019a] invokes AllReduce many times within each model pass to exchange activations. In ML software stacks, collective communication libraries (CCLs) like NCCL implement algorithms for collective primitives, interfacing with ML application libraries (e.g., PyTorch [PyTorch Contributors, 2024], TensorFlow [TensorFlow Developers, 2024]) at one end and network transport protocols at the other. For AllReduce, NCCL chooses between multiple algorithms based on the number of GPUs and the amount of data to be reduced. Since buffer sizes in modern ML workloads tend to be large [Shah et al., 2023], CCLs leverage bandwidth-optimal algorithms (e.g., Ring) to minimize the AllReduce time.
Most CCLs, including NCCL, adopt a bulk-synchronous model where all GPUs synchronize before the collective, guaranteeing data correctness and enabling efficient parallel communication algorithms that rely on all ranks to implement the collective concurrently. However, performance degrades significantly when one or more GPUs are slow to reach the synchronization point. The slowest GPU (the straggler) dictates overall performance due to tail effects [Warraich et al., 2025, Dean and Barroso, 2013, Wang et al., 2024a, Gangidi et al., 2024, Li et al., 2014]. Even with multiple slower GPUs, there is typically only a single straggler since the probability of multiple GPUs completing exactly simultaneously is extremely small. Straggler delays may stem from hardware issues like thermal throttling and power supply, or runtime factors like network congestion, background noise, and unavoidable skewed compute allocations across GPUs [Wu et al., 2024, Jiang et al., 2024, Grattafiori et al., 2024, Xiong et al., 2024]. Recent work has highlighted the significant impact of stragglers on datacenter-scale ML jobs [Jiang et al., 2024, Lin et al., 2025], and our experiments fine-tuning Llama-3.2 (1B/3B) across different hardware setups (Fig. 2(a)) reveal straggler effects of up to 30 ms even within individual multi-GPU servers.
Method Domain Backend Application Effect on Loss/Convergence
[Harlap et al., 2016] Scale-out Custom C++ runtime DP No
[Karakus et al., 2017] Scale-out MPI DP Yes
[Wang et al., 2020] Scale-out NCCL DP, TP No
[Won et al., 2024] Irregular scale-up Simulation DP, TP No
[Zhao et al., 2024a] Scale-out Custom CUDA runtime DP, TP, EP No
[Warraich et al., 2025] Scale-out Gloo & custom transport DP Yes
StragglAR (Ours) Scale-up NCCL DP, TP No
Table 3: StragglAR vs. prior work (DP=data parallel, TP=tensor parallel, EP=expert parallel). While StragglAR targets the scale-up domain, it also applies to switched homogeneous scale-out domains, like in today’s rail-optimized topologies for ML.
Straggler mitigation. Some prior works attempt to identify and remove straggling devices or thoroughly investigate root causes [Jiang et al., 2024, Lin et al., 2025]. However, removing stragglers wastes compute capacity, and only applies to severe stragglers rather than the routine compute heterogeneity we observe regularly, even at small scales. Moreover, stragglers arise from many sources, making diagnosis difficult and incomplete [Lin et al., 2025]. Recent works obviate straggler delays by approximating or dropping the straggler’s data during training [Warraich et al., 2025, Harlap et al., 2016, Karakus et al., 2017, Recht et al., 2011]. While these approximations and asynchronous techniques can progress without waiting for stragglers, they are limited to data-parallel training and can impact model convergence. In contrast, StragglAR is a fundamentally new AllReduce algorithm that ensures accurate, unmodified reductions and is agnostic to the AllReduce use case, making it applicable to both data-parallel training and tensor-parallel training/inference. Like all collective algorithms, StragglAR requires participation from all GPUs (including the straggler), but accelerates the collective itself. Other works mitigate stragglers in datacenter-scale ML jobs through systems-level workload rebalancing and runtime strategies [Wu et al., 2024, Zhao et al., 2024a], improving efficiency by adapting task placement or dynamically selecting among different known collective algorithms based on profiling. For example, our investigation of AdapCC [Zhao et al., 2024a] finds that it relies classical Tree algorithms, which sacrifice bandwidth to improve latency. In contrast, StragglAR is not a runtime controller, but a new bandwidth-efficient parallel algorithm for AllReduce that changes the communication schedule itself to provably reduce communication complexity. It can be integrated with different runtime systems that implement these algorithms.
AllReduce algorithms. Collective algorithms are designed based on knowledge of the underlying network topology. Inter-GPU networks typically fall into two categories: scale-up (homogeneous high-bandwidth links within a node or rack) and scale-out (heterogeneous or unpredictable bandwidth across multiple nodes or racks). Recent works have synthesized new AllReduce algorithms for heterogeneous networks with asymmetric bandwidth [Shah et al., 2023, Zhao et al., 2024a, Wang et al., 2020, Won et al., 2024, Zhang et al., 2024]. However, for the homogeneous networks that define modern scale-up domains, these synthesizers invariably converge to bandwidth-optimal classical algorithms, such as Ring and Recursive Halving/Doubling [Shah et al., 2023, Wang et al., 2020]. Consequently, state-of-the-art CCLs like NCCL continue to implement classical algorithms for collective communication in scale-up domains [Hu et al., 2025]. We provide Table 3 to concretely situate our contribution in the space of collective algorithms and straggler mitigation strategies. We note that while we focus on the scale-up domain for StragglAR, it applies to any homogeneous switched network, including rail-optimized scale-out domains [NVIDIA, 2022].
@@ Appendix A Extended Related Work
Optimal collective communication. Over decades, researchers have implemented collective communication primitives with fast algorithms that are carefully attuned to latency and bandwidth trade-offs. With today’s large ML models that require communicating large amounts of data, bandwidth-efficient algorithms are of primary focus [Narayanan et al., 2021, Fei et al., 2021]. In homogeneous networks, well-known parallel algorithms, such as Ring, recursive halving/doubling, etc., enable fast, bandwidth-efficient AllReduce by heavily parallelizing communication in every round of the algorithm [Patarasuk and Yuan, 2009, Rabenseifner, 2004, Rabenseifner and Träff, 2004, Thakur et al., 2005, Bruck et al., 1994, Sanders et al., 2009]. For heterogeneous networks, recent works synthesize custom collective algorithms by using the hardware topology or parallelization strategy as an input [Shah et al., 2023, Kim et al., 2024, Cai et al., 2021, Zhao et al., 2024b, Laskar et al., 2024, Won et al., 2024, Mahajan et al., 2023]; optimizing a collective algorithm to minimize the latency and bandwidth cost, however, is a challenging combinatorial optimization problem that often requires hand-tuned heuristics to constrain the search space [Shah et al., 2023]. Further, some approaches optimize AllReduce for a specific topology [Jain and Sabharwal, 2010, De Sensi et al., 2024, Wang et al., 2023a]. Our work extends this line of research by optimizing AllReduce in settings with a straggler for distributed ML, where bandwidth-efficient algorithms are required due to the large buffer sizes [Narayanan et al., 2021, Shah et al., 2023, NVIDIA, 2024a].
@@
Rebalancing workloads. Recent works improve the efficiency of distributed ML by optimizing the model parallelization strategy, which can affect data/work sharding across GPUs and the collective communication incurred [Jia et al., 2019, Zhang et al., 2024, Zheng et al., 2022, Shoeybi et al., 2019b, Rajbhandari et al., 2020, Grattafiori et al., 2024, Narayanan et al., 2021, Unger et al., 2022, Zhao et al., 2024a]. Vanilla parallelization strategies evenly shard or replicate the compute and memory load (i.e., model weights, input batches, activations, optimizer states, etc.) across GPUs within a group to achieve speedups from parallelization. This balances the computation and communication load experienced by each worker, but could introduce resource idling when there are slow GPUs in the group. Our experiments in Fig. 7 and §B highlight significant straggler effects even with uniform workload placement. Other works identify severe stragglers and balance the workload dynamically to limit their impact [Harlap et al., 2016, Wu et al., 2024, Wang et al., 2023b, Lin et al., 2025]. StragglAR directly addresses stragglers at a lower layer of abstraction, within the AllReduce kernel, and thus expands on the problems tackled by these works.
@@
Sending less data. Various methods leverage the resilience of stochastic optimization algorithms used in training to reduce the communication volume by selectively communicating data with high statistical value [Warraich et al., 2025, Li and Hoefler, 2022, Wang et al., 2023c], synchronizing gradients with varying frequencies [Giladi et al., 2023, Wang and Varma, 2024, Liu et al., 2022, Karakus et al., 2017, Yang et al., 2020, 2020], or sending lower-precision or compressed data [Alistarh et al., 2017, Bernstein et al., 2018, Wang et al., 2022a, 2024b, 2023c, Lu et al., 2022, Tang et al., 2021, Renggli et al., 2019, M Abdelmoniem et al., 2021, Wang et al., 2023d]. StragglAR is orthogonal to these works because it optimizes the collective algorithm itself (i.e., exactly how the data is transmitted as opposed to how much) and can operate given any buffer size, including after compression. Further, because StragglAR directly modifies the collective algorithm, as opposed to the data that is sent or the frequency of synchronization, it also supports AllReduce for tensor parallelism, which is used in both training and inference [Shoeybi et al., 2019a].
Compute-communication overlap. Recent works propose fine-grained overlap of compute and communication to mitigate synchronization delays and reduce communication overheads in distributed ML [Ismayilov et al., 2023, Jangda et al., 2022, Wang et al., 2023e, NVIDIA, 2024c, Punniyamurthy et al., 2024, Liang et al., 2024, Pati et al., 2024, Chen et al., 2024, Wang et al., 2022b]. Although this approach may be useful for small, irregular communication patterns, such as expert-parallel AllToAll [Liu et al., 2024], it is less useful with large data buffers that require aggregation and are bottlenecked by network bandwidth [Dryden et al., 2018, Patarasuk and Yuan, 2009]. Thus, bulk data transfers, like gradient synchronization in data-parallel training or activation aggregation in tensor-parallel training with large batch sizes, rely on bandwidth-optimal collective communication. Further, these works consider overlapping compute and communication at higher layers of abstraction, but rely on the same algorithms to implement the collective communication. StragglAR instead brings the idea of overlap to the communication algorithm itself, and uses this to directly address the straggler problem. Other works propose entirely abandoning synchronization barriers and relying on resilience of stochastic gradient descent to enable convergence during training [Nabli et al., 2023, Tyurin et al., 2024, Su et al., 2022, Tyurin and Richtarik, 2024, De Vos et al., 2023, Recht et al., 2011]. However, these approaches impact model convergence, limiting applicability, and do not generalize to tensor-parallel training and inference, which require accurate and complete reductions.
Appendix B Persistent Stragglers
While recent works [Jiang et al., 2024, Lin et al., 2025] have shown persistent stragglers in ML jobs with thousands of GPUs, we show that even smaller multi-GPU jobs encounter such stragglers. We run Llama-3.2 fine-tuning jobs in three hardware environments: (1) 4 A100s with 40GB memory and NVLink 3.0 fully-connected mesh network in the Perlmutter supercomputer [NERSC, 2025], (2) 4 A100s with 80GB memory in RunPod [RunPod, Inc., 2025], and (3) 8 A100s with 80GB memory in RunPod. Both RunPod servers use NVLink 3.0 with an NVSwitched any-to-any network. For every environment, we repeat the same workload three times.
We encounter stragglers in all nine runs. We define the straggler delay as $T_{delay}=T_{rest}-T_{straggler}$, where $T_{straggler}$ is the AllReduce execution time of the straggler — the GPU that started the AllReduce last (rank 0 in Fig. 1) — and $T_{rest}$ is the AllReduce execution time of the second-slowest rank. Note that $T_{straggler}$ is the shortest AllReduce time among all GPUs because the straggler spends no time waiting for others before communicating, and our measure of $T_{delay}$ is conservative since it only encodes the idle time of the second-slowest GPU. We find a median straggler delay of more than 8 ms and a mean delay of 9.45 ms. The straggler GPU rank persists not only across iterations, but also across runs. Fig. 7 shows that one specific GPU is the slowest in a communication group in 98% of iterations, and is also the persistent straggler over multiple independent runs of the same workload on the same VM. Fig. 7(a) shows these findings also generalize to multiple stragglers. Our experiments document (1) the severity of straggler GPUs that consistently delay AllReduce, and (2) that the same GPU rank is typically the straggler, allowing us to use offline profiling to bootstrap StragglAR. These findings also suggest that routine stragglers may even appear from intrinsic and unavoidable GPU hardware heterogeneity (even with the same GPU model).
(a) Perlmutter, 4 A100 SXM, Llama-3.2-1B, batch_size=32.
(b) RunPod, 4 A100 SXM, Llama-3.2-3B, batch_size=32.
(c) RunPod, 8 A100 SXM, Llama-3.2-3B, batch_size=64.
Appendix C Proof of Communication Complexity